OpenStudy (anonymous):

Choose the correct simplification of the expression (2x2y6z5)(5x4y5z3). 10 x8y30z15 7 x6y11z8 7x8y30z15 10x6y11z8

OpenStudy (anonymous):

just like what u did at the last one

OpenStudy (anonymous):

\[x^2.x^4=x^{2+4}=x^6\]

OpenStudy (anonymous):

can u continue?

OpenStudy (mertsj):

Multiply the coefficients and add the exponents of the factors that hae the same base.

OpenStudy (mertsj):

Time for you to learn the laws of exponents, sweetheart.

OpenStudy (anonymous):

lmao ^_^

OpenStudy (anonymous):

i know.